The equation h=-16^2 +112t gives the height (h) of an arrow shot upward from the ground with an initial velocity of 112ft/s, where t represents the time after the arrow leaves the ground. Using that equation, find the height of the arrow after (a)2seconds (b)5 seconds
Answer by Alan3354(69443)

It's h(t) = -16t^2 +112t
Sub 2 for t
h(2) = -16*4 + 112*2 = 224-64
h(2) = 160 feet
-------------
Do the same with 5